### v1

@@ 7.2. Expanding the Perceived Boundary of Documentable Knowledge
Our study revealed a significant gap between the knowledge that lab members possessed and what they felt was appropriate to document. Many participants hesitated to share context-specific experiences that may be generalizable and beneficial to their peers. This reluctance appears to stem from an implicit expectation that documentation should be universal, authoritative, and permanent—creating a high bar for contribution that excludes valuable experiential knowledge and partial solutions. As an alternative, participants also recognized significant value in case-based information and partial policies that could provide personalized guidance even when comprehensive documentation was unavailable. This challenge is well-studied: from early work on organizational memory systems (Conklin, 1997) to research on knowledge sharing barriers (Ackerman et al., 2013), studies have consistently identified the difficulty of capturing informal knowledge due to cultural and technical barriers toward outcome-focused documentation. As a result, much valuable knowledge remains trapped in individual experience rather than being shared with future lab members.
To address this limitation, we propose a more expansive model of organizational memory that embraces contextual knowledge alongside formal policies. Future systems should support lightweight contribution mechanisms that allow members to share experiences, partial solutions, and case-based knowledge without the burden of creating universally applicable documentation. Furthermore, systems can provide guided pathways that help users navigate between formal documentation and informal human expertise when contributing knowledge. This concept aligns with early systems like Answer Garden (Ackerman and Malone, 1990). In addition, establishing common ground between lab directors and students, and fostering a welcoming environment even when the knowledge is canonical, may encourage students to contribute—an aspect that was not captured in our field study.
@@ 7.5. Limitations
Our study has several limitations that may impact the generalizability of the findings. First, our field deployment involved a relatively small sample of participants (21 total) across four university research labs, all from STEM fields with high technical literacy. This limits generalizability to organizations with different technical backgrounds, hierarchical structures, or cultural norms around knowledge sharing. Second, the one-month deployment period was sufficient for observing initial adoption patterns but insufficient for evaluating the longer-term effectiveness of organizational memory systems. Our study focused on discovering potential effects rather than proving sustained efficacy. Third, conducting the study during the summer when lab activities were less active than usual may not reflect typical year-round organizational memory practices.

### v2

@@ 7.2. Expanding the Perceived Boundary of Documentable Knowledge
Our study revealed a gap between the knowledge that lab members possessed and what they felt was appropriate to document. Many participants hesitated to share context-specific experiences that may be generalizable and beneficial to their peers. This reluctance appears to stem from an implicit expectation that documentation should be universal, authoritative, and permanent—creating a high bar for contribution that excludes valuable experiential knowledge and partial solutions. Participants also recognized value in case-based information and partial policies that could provide personalized guidance when comprehensive documentation was unavailable. This challenge is well-studied: from early work on organizational memory systems (Conklin, 1997) to research on knowledge sharing barriers (Ackerman et al., 2013), studies have consistently identified the difficulty of capturing informal knowledge due to cultural and technical barriers toward outcome-focused documentation. Prior work on enterprise wikis found that members only occasionally contributed despite recognizing documentation value, often requiring explicit requests from managers (Munson, 2008; Kiniti and Standing, 2013). CHOIR addressed this barrier by embedding documentation opportunities within natural conversation flows. Participants noted that the knowledge extraction feature made contributing informal knowledge feel less burdensome, as it transformed ongoing discussions into documentation without requiring separate writing effort. This suggests that conversational AI systems can provide lightweight pathways for capturing experiential knowledge that traditional documentation systems have struggled to elicit.
Building on these findings, we propose a more expansive model of organizational memory that embraces contextual knowledge alongside formal policies. Future systems should support lightweight contribution mechanisms that allow members to share experiences, partial solutions, and case-based knowledge without the burden of creating universally applicable documentation. Furthermore, systems can provide guided pathways that help users navigate between formal documentation and informal human expertise when contributing knowledge. This concept aligns with early systems like Answer Garden (Ackerman and Malone, 1990). In addition, establishing common ground between lab directors and students, and fostering a welcoming environment even when the knowledge is canonical, may encourage students to contribute.
@@ 7.3. AI Agents as Mediator in Organizations
From a socio-technical perspective, CHOIR’s mediating role can be understood through the lens of boundary objects (Star and Griesemer, 1989). Organizational memory documents serve as boundary objects with interpretive flexibility, where directors view them as comprehensive repositories of lab knowledge that they want to convey to students, while students see them as practical guides for navigating their immediate questions and expectations. In our work, an AI agent mediates existing boundary objects to each group, and the AI agent itself is a new boundary object that each group can access and communicate with. Extending recent work on AI systems as boundary objects in creative and participatory contexts (Chung and Adar, 2023; Ayobi et al., 2021; Porquet et al., 2025; Guridi et al., 2025), CHOIR shows how AI can serve this role in organizational memory, dynamically adapting the same knowledge base to serve different communities’ needs. This mediation extends organizational knowledge from context-dependent exchanges to persistent and reusable organizational memory that different members can retrieve and apply to their own situations.
@@ 7.5. Limitations
Our study has several limitations that may impact the generalizability. First, our field deployment involved a small sample of participants (22 total) across four university research labs, all from STEM fields with high technical literacy. This limits generalizability to organizations with different technical backgrounds, hierarchical structures, or cultural norms around knowledge sharing. Second, the one-month deployment period was sufficient for observing initial adoption patterns but insufficient for evaluating the longer-term effectiveness of organizational memory systems. Our study focused on discovering potential effects rather than proving sustained efficacy.
Third, because the study was conducted during the summer, the findings primarily reflect the onboarding experiences of new remote students. Although this context provided insights into early-stage lab integration, usage patterns during regular academic semesters may differ. We plan to conduct a long-term follow-up with participating labs to examine if and how usage and perceptions evolve across different academic periods. Additionally, the study setup required preparing initial documentation for participating labs by creating new seed documents or converting existing materials into standardized formats. We acknowledge that these prepared documents may have affected ecological validity compared with naturally evolved documentation.
Fourth, one participating lab (L4) was led by the corresponding author, which may introduce potential bias despite our efforts to minimize researcher influence. While the corresponding author limited their role to managing the process, answering questions, and approving documents, and the first and second authors did not participate as study subjects, peer-pressure inherent in being in the same lab or the natural interest in a research project that is thematically similar to their own research could have influenced the L4 participants’ engagement patterns or feedback. However, the key findings and discussion points are valid as they are not disproportionately driven by L4 data alone.
Lastly, our study lacks a controlled baseline and quantitative performance measures beyond descriptive statistics, which limits our ability to make causal claims about CHOIR’s effectiveness compared to alternatives such as wiki pages or existing lab practices (e.g., Google Docs/Slack Pinned Messages). As a result, our findings are descriptive rather than comparative, and further research is needed to determine how specific features of CHOIR contribute to its impact.
8. Conclusion
In this paper, we presented CHOIR, a chatbot-mediated organizational memory system that integrates document-grounded Q&A, knowledge sharing, conversation-based knowledge extraction, and AI-assisted document updates within workplace communication platforms. Through a formative study with 15 participants, we identified key challenges in managing organizational memory in research labs, which guided our design of CHOIR. We then deployed CHOIR across four university research labs for one month, observing that it supported knowledge retrieval and documentation practices, with 42% of questions answered from existing documentation and 1,994 words added to organizational memory. Our findings revealed a fundamental privacy-awareness tension in AI-mediated organizational memory systems, where students’ privacy-seeking behaviors, driven by concerns about social evaluation, limited directors’ awareness of knowledge gaps. We found that conversational agents can serve as facilitators rather than replacements for human communication, offering multiple channels for more effective human communication while identifying gaps in organizational documentation. Our work provides empirical observations and design implications for integrating AI-mediated organizational memory into workplace communication platforms, highlighting the importance of addressing psychological safety, power dynamics, and cultural norms in hierarchical organizations. These findings offer insights for designing knowledge management systems that balance individual privacy with organizational transparency and support capturing contextual knowledge that often remains undocumented.
Appendix A System Prompts
This appendix provides the complete system prompts used in CHOIR for reproducibility. Model configuration details are provided in Section 4.5.2.
A.1. Question Answering Prompt
The following prompt is used for document-grounded Q&A (Feature 1). User names are replaced with pseudonyms before being sent to the API for privacy protection.
You are CHOIR, a helpful AI assistant for [Organization Name].
Think of yourself as a knowledgeable senior student or friendly
professor who’s always ready to help with questions.
I have access to the organization’s documentation and knowledge
base, so I can help you find information and provide guidance
based on what we have documented.
Organization Information:
- Organization: [Organization Name]
- About: [Organization Description]
- Today’s date: [Current Date]
- Workspace: [Workspace Name]
IMPORTANT: First, determine if you can answer the user’s question
based ONLY on the provided references. Do NOT use your general
knowledge or make assumptions beyond what’s explicitly stated in
the references.
Then respond with a JSON object containing:
1. "canAnswer": true/false - whether the references contain
   sufficient information to answer the question
2. "response": your answer based on the references, or a friendly
   explanation that you couldn’t find the information
Guidelines for answering:
- Be friendly, approachable, and helpful - like a senior colleague
  who genuinely wants to help
- Answer ONLY based on the provided references below - do NOT add
  general knowledge
- If multiple references contain conflicting information,
  prioritize the first reference in the list
- If you cannot answer based on the references, encourage the user
  to ask others or start a discussion to help improve our
  documentation
- When you cannot fully answer a question, mention which related
  references you found and what information they contain
- Use a warm, academic tone - professional but not overly formal
==== REFERENCES ====
[Retrieved document chunks]
==== CONVERSATION HISTORY ====
[Previous messages in the thread]
==== CURRENT QUESTION ====
[User’s question]
Analyze whether you can answer based on the documentation and
provide your response as JSON.
A.2. Knowledge Extraction Prompt
The following prompt is used to extract documentable knowledge from conversations (Feature 3). User names are replaced with pseudonyms before being sent to the API.
You are CHOIR, a documentation specialist. Extract organizational
knowledge from the conversation that would be valuable for future
reference.
Only extract information that establishes policies, procedures, or
reusable knowledge for the organization. Do NOT extract personal
preferences, individual decisions, or casual conversation.
Start with a descriptive markdown section title (# [Topic Name]),
then write the information in natural paragraph format. Only include
facts that are directly stated in the conversation - do not add
explanations, interpretations, or implications.
Do not include personal conversation details like who said what.
Always preserve any URLs mentioned.
Conversation:
[Conversation messages]
What information is shared in the conversation that should be
documented?
A.3. Document Update Prompt
The following prompt is used to integrate extracted knowledge into existing documentation (Feature 4).
You are a document editor. Improve existing content by integrating
the provided knowledge.
Existing content type: [paragraph or list]
Rules:
- PRIORITIZE updating/replacing existing content when knowledge
  provides better, more accurate, or more comprehensive information
- If knowledge contradicts existing content, prefer the knowledge
  (assume it’s more current/accurate)
- If knowledge complements existing content without contradiction,
  add it in matching format
- If knowledge provides more specific details about existing points,
  merge them into improved versions
- Preserve all URLs from the knowledge
- Use single-level lists only (no nested bullets)
- No headings or section titles
- Return original only if knowledge adds no meaningful value
Approach: Update first, then add if needed. Create the most accurate
and comprehensive version.
Wrap your response in <markdown> tags.
File: [File Name] - Section: [Section Name]
Existing content:
<markdown>
[Current section content]
</markdown>
Knowledge to integrate:
[Extracted knowledge]